# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ B.2 Proof of Theorem 2
$=$ $\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}+a_{i}/\sqrt{T}|-\mathopen{}\mathclose{{}\left(\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}|}\right)$
@@
for all $\mathbf{a}$. Note that the inequality (B.2) follows from the definition of $\widehat{\boldsymbol{\beta}}$. Thus, we see that
@@
By the Argmin Theorem (Geyer, 1996), we can claim that $\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V(\mathbf{a})$, which implies that $\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V(\mathbf{a})$, which would prove the Theorem. In what follows we show that $V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptV(\mathbf{a})$ for all $\mathbf{a}$. Note that
@@
As $T\to\infty$ we have $\mathbf{a}^{\prime}\mathopen{}\mathclose{{}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbf{w}_{t}\mathbf{w}_{t}^{\prime}}\right)\mathbf{a}\to\mathbf{a}^{\prime}\boldsymbol{C}_{w}\mathbf{a}$. Note that $\{v_{t}y_{t-l}\},l\geq 1$, has mean 00, autocovariance function $\gamma(\cdot)$ such that $\sum_{h=-\infty}^{\infty}|\gamma(h)|<\infty$, and autocorrelation coefficient $\phi_{l}$ such that $\sum_{j=0}^{\infty}\phi_{j}\neq 0$. Thus, we can apply the CLT under weak dependence (see  Billingsley, 1995, Thm. 27.4) to obtain
[…]
@@
[…]
@@
[…]
@@
which shows that $V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptV(\mathbf{a}).$blacksquare
Remark B.2:
Under the common AR($p$) restriction (see Remark 1), $v_{t}=\omega_{t}$ and $E(v_{t}|y_{t-l-1},y_{t-l-2},\dots)=0$, $\forall l\geq 1$. Thus, if $\lambda\rightarrow 0$ and $T^{\frac{1-c}{2}}\lambda\rightarrow\infty$, $c\in[0,1)$, then Theorem 2 holds with $\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}(V(\mathbf{a}))=\mathbf{C}_{w}^{-1}\mathbf{m}\sim N(\bf 0_{n+p_{y}},\sigma_{v}^{2}\mathbf{C}_{w})$ (Fu and Knight (2000), Thm. 2) and Theorem 3 ensures $P\mathopen{}\mathclose{{}\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}\right)=1-o(e^{-T^{c}})$ for $c\in[0,1)$ (Zhao and Yu (2006), Thm. 1).
B.3 Proof of Theorem 3
@@
.$\mathcal{E}.1_{T}$ $=$ $\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|\widehat{\mathbf{c}}_{i}(11)^{-1}_{i}b_{i}(1)}\right|<\sqrt{T}\mathopen{}\mathclose{{}\left(|\beta_{i}^{*}|-\frac{\lambda}{2T}\mathopen{}\mathclose{{}\left|\widehat{\mathbf{c}}_{i}(11)^{-1}Sign(\beta_{i}^{*})}\right|}\right)}\right\},\ i=1,\dots,s,n+1,\dots,n+s_{y}$
.$\mathcal{E}.2_{T}$ $=$ $\mathopen{}\mathclose{{}\left\{|b_{i}-b_{i}(2)|\leq\frac{\lambda\varphi}{2\sqrt{T}}}\right\},\ i=s+1,\dots,n,n+s_{y}+1,\dots,n+p_{y}\ \ ,$
where $\widehat{\mathbf{c}}_{i}(11)_{i}$, $b_{i}$, $b_{i}(1)$ and $b_{i}(2)$ are elements of $\widehat{\mathbf{C}}_{w}(11),\ \mathbf{b}=\mathopen{}\mathclose{{}\left(\widehat{\mathbf{C}}_{21}(\widehat{\mathbf{C}}_{11})^{-1}\mathbf{W}(1)\mathbf{v}}\right),\ \mathbf{b}(1)=\frac{1}{\sqrt{T}}\mathbf{W}(1)\mathbf{v}$ and $\mathbf{b}(2)=\sqrt{T}\mathbf{W}(2)\mathbf{v}$, respectively. .$\mathcal{E}.1_{T}$ implies that the signs of the relevant predictors are correctly estimated, while .$\mathcal{E}.1_{T}$ and .$\mathcal{E}.2_{T}$ together imply that the signs of the irrelevant predictors are shrunk to zero. To show $P\mathopen{}\mathclose{{}\left(\exists\lambda\geq 0:Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}\right)\rightarrow 1$, it is sufficient to show that $P\mathopen{}\mathclose{{}\left(\exists\lambda\geq 0:Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}\right)\geq P\mathopen{}\mathclose{{}\left(\mathcal{E}.1_{T}\cap\mathcal{E}.2_{T}}\right)$ (see Proposition 1 in  Zhao and Yu, 2006). Using the identity of $1-P\mathopen{}\mathclose{{}\left(\mathcal{E}.1_{T}\cap\mathcal{E}.2_{T}}\right)\leq P\mathopen{}\mathclose{{}\left(\mathcal{E}.1_{T}^{C}}\right)+P\mathopen{}\mathclose{{}\left(\mathcal{E}.2_{T}^{C}}\right)$ we have that
$P\mathopen{}\mathclose{{}\left(\mathcal{E}.1_{T}^{C}}\right)+P\mathopen{}\mathclose{{}\left(\mathcal{E}.2_{T}^{C}}\right)$$\leq$ $\sum_{i=1}^{s,n+1,\dots,n+s_{y}}P\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{T}}|\widehat{\mathbf{c}}_{i}(11)^{-1}\mathbf{w}_{i}^{\prime}\mathbf{v}|\geq\sqrt{T}\mathopen{}\mathclose{{}\left(|\beta_{i}^{*}|-\frac{\lambda}{2T}|\widehat{\mathbf{c}}_{i}(11)^{-1}Sign(\beta_{i}^{*})|}\right)}\right)$
$+$ $\sum_{i=1}^{s+1,\dots,n,n+s_{y}+1,\dots,n+p_{y}}P\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{T}}|b_{i}-\mathbf{w}_{i}^{\prime}\mathbf{v}|\geq\frac{\lambda\varphi}{2\sqrt{T}}}\right)=I_{T}+II_{T}\ \ .$
Note that by Assumption 5, $\widehat{\psi}_{max}^{w}\geq\widehat{\psi}_{min}^{w}\geq 0$, hence
@@
for some $c_{0}>0$ (see  Zhao and Yu, 2006, Thm. 3 and 4). Therefore, by the union bound, Markov’s inequality and the mixingale concentration inequality (see Hansen, 1991, Lemma 2), we have that
@@
where $c_{1}>2$ (see Assumption 4 in the main text). Conducting a similar analysis for $II_{T}$, and considering that by assumption $\sqrt{T}\lambda\rightarrow\lambda_{0}\geq 0$, we obtain $P\mathopen{}\mathclose{{}\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}\right)\rightarrow 1.$blacksquare
B.4 Proof of Theorem 4
Before providing the proof of Theorem 4, we introduce some important definitions.
Definition B.1:
@@ B.4 Proof of Theorem 4
Definition B.2:
(Mixingale, Davidson (1994), ch. 16). The sequence of pairs $\mathopen{}\mathclose{{}\left\{X_{t},\mathcal{F}}\right\}_{-\infty}^{+\infty}$ in a filtered probability space $\mathopen{}\mathclose{{}\left(\Omega,\mathcal{F},P}\right)$ where the $X_{t}$ are integrable r.v.s is called $L_{p}$-mixingale if, for $p\geq 1$, there exist sequences of non-negative constants $\mathopen{}\mathclose{{}\left\{d_{t}}\right\}_{-\infty}^{+\infty}$ and $\mathopen{}\mathclose{{}\left\{\nu_{m}}\right\}_{0}^{\infty}$ such that $\nu_{m}\rightarrow 0$ as $m\rightarrow\infty$ and
@@
Definition B.3:
(Near-Epoch Dependence, Davidson (1994), ch. 17). For a possibly vector-valued stochastic sequence $\mathopen{}\mathclose{{}\left\{\mathbf{V}_{t}}\right\}_{-\infty}^{+\infty}$, in a probability space $\mathopen{}\mathclose{{}\left(\Omega,\mathcal{F},P}\right)$ let $\mathcal{F}_{t-m}^{t+m}=\mathbf{\sigma\mathopen{}\mathclose{{}\left(\mathbf{V}_{t-m},\dots,\mathbf{V}_{t+m}}\right)}$, such that $\mathopen{}\mathclose{{}\left\{\mathcal{F}_{t-m}^{t+m}}\right\}_{m=0}^{\infty}$ is a non-decreasing sequence of $\mathbf{\sigma}$-fields. If for $p>0$ a sequence of integrable r.v.s $\mathopen{}\mathclose{{}\left\{X_{t}}\right\}_{-\infty}^{+\infty}$ satisfies
@@
To simplify the analysis, we frequently make use of arbitrary positive finite constants $C$, as well as of its sub-indexed version $C_{i}$, whose values may change from line to line throughout the paper, but they are always independent of the time and cross-sectional dimension. Generic sequences converging to zero as $T\rightarrow\infty$ are denoted by $\zeta_{T}$. We say a sequence $\zeta_{T}$ is of size $-\phi_{0}$ if $\zeta_{T}=O(T-\phi_{0}-\varepsilon)$ for some $\varepsilon>0$.
Remark B.2:
@@
Under Assumption 1 of the main text the process $\{x_{i,t}\}$ is $L_{2b_{1}}$-NED of size $-a$, with $a\geq 1$, while the process $\{q_{i,t}\}$ is $L_{2c_{1}}$-NED of size $-d$, with $d\geq 1$. By Theorems 17.5 in ch.17 of Davidson (1994), they are also $L_{b_{1}}$ and $L_{c_{1}}$-Mixingale, respectively. In later theorems, the NED order and sequence size are important for asymptotic rates. Assumption 4 requires $\mathbf{q}_{t}$ to have slightly more moments than $c_{1}$. More moments mean tighter error bounds and weaker tuning parameter conditions, but a high $c_{2}$ imposes stronger model restrictions. Under strong dependence, fewer moments are needed, and the reduction from $c_{2}$ to $c_{1}$ reflects the cost of allowing greater dependence through a smaller mixing rate.
The proof of Theorem 4 follows that of Theorem 1 in Adamek et al. (2023). We need of the two following auxiliary Lemmas.
Lemma B.5:
Let $\mathcal{B}_{T}\coloneqq\mathopen{}\mathclose{{}\left\{\underset{i\leq n_{T}}{\text{max}}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\widehat{\boldsymbol{\gamma}}_{i}-\boldsymbol{\gamma}_{i}}\right|}\right|_{\infty}\leq C}\right\}$. Under Assumptions 1 and 3, for $a\geq 1$ and a sequence $\zeta_{T}\rightarrow 0$ such that $\zeta_{T}\leq\frac{n_{T}p_{i}}{e}$, if $\sqrt{T}\geq C(n_{T}p_{i})^{\frac{b_{1}+a-1}{a(b_{1}-1)}}\zeta_{T}^{-\frac{b_{1}+a-1}{a(b_{1}-1)}}$, then $\Pr(\mathcal{B}_{T})\geq 1-3\zeta_{T}\rightarrow 1$.
Proof of Lemma B.5: By the union bound we have
$\Pr(\mathcal{B}_{T}^{c})$$\leq$ $\sum_{i=1}^{n_{T}}\sum_{l=1}^{p_{i}}\Pr\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left|\sum_{t=1}^{T}(x_{i,t}x_{i,t-l}-E(x_{i,t}x_{i,t-l}))}\right|>CT}\right)$

### v3

@@ A.3 Proof of Theorem 1
$=$ $\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}+a_{i}/\sqrt{T}|$
$-\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}|}}\right)$
@@
for all $\mathbf{a}$. Note that the inequality (A.3) follows from the definition of $\widehat{\boldsymbol{\beta}}$. Thus, we see that
@@
By the Argmin Theorem (geyer1996), we can claim that $\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V(\mathbf{a})$, which implies that $\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V(\mathbf{a})$, which would prove the Theorem. In what follows we show that $V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptV(\mathbf{a})$ for all $\mathbf{a}$. Note that
@@
As $T\to\infty$ we have $\mathbf{a}^{\prime}\mathopen{}\mathclose{{\left(\frac{1}{T}\sum_{t=1}^{T}\mathbf{w}_{t}\mathbf{w}_{t}^{\prime}}}\right)\mathbf{a}\to\mathbf{a}^{\prime}\boldsymbol{C}_{w}\mathbf{a}$. Note that $\{v_{t}y_{t-l}\},l\geq 1$, has mean 0, autocovariance function $\gamma(\cdot)$ such that $\sum_{h=-\infty}^{\infty}|\gamma(h)|<\infty$, and autocorrelation coefficient $\phi_{l}$ such that $\sum_{j=0}^{\infty}\phi_{j}\neq 0$. Thus, we can apply the CLT under weak dependence (see billingsley1995, Thm. 27.4) to obtain
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}v_{t}\mathbf{a}^{\prime}\mathbf{w}_{t}\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptN\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n}^{\prime},\boldsymbol{\mu}_{vy}}}\right)^{\prime},\mathbf{a}^{\prime}\mathopen{}\mathclose{{\left({\begin{array}[]{cc}\sigma_{v}^{2}\mathbf{C}_{u}&\boldsymbol{0}_{n\times p_{y}}\\ \boldsymbol{0}_{p_{y}\times n}&\boldsymbol{\Gamma}_{vy}\\ \end{array}}}}\right)\mathbf{a}}}\right)\ \ .$
@@
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}v_{t}\mathbf{a}^{\prime}\mathbf{w}_{t}\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\mathbf{a}^{\prime}\mathbf{m}\ \ ,$
@@
Applying Slutsky’s theorem, we have $I(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\mathbf{a}^{\prime}\mathbf{C}_{w}\mathbf{a}-2\mathbf{a}^{\prime}\mathbf{m}.$
@@
which shows that $V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptV(\mathbf{a}).$ $\square$
Remark 2:
Under the common AR($p$) restriction (see Remark 2 of the main text), $v_{t}=\omega_{t}$ and $E(v_{t}|y_{t-l-1},y_{t-l-2},\dots)=0$, $\forall l\geq 1$. Thus, if $\lambda\rightarrow 0$ and $T^{\frac{1-c}{2}}\lambda\rightarrow\infty$, $c\in[0,1)$, then Theorem 1 holds with $\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}(V(\mathbf{a}))=\mathbf{C}_{w}^{-1}\mathbf{m}\sim N(\bf 0_{n+p_{y}},\sigma_{v}^{2}\mathbf{C}_{w})$ (Fu2000, Thm. 2) and Theorem 2 ensures $P\mathopen{}\mathclose{{\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)=1-o(e^{-T^{c}})$ for $c\in[0,1)$ (zhaoyu2006, Thm. 1).
A.4 Proof of Theorem 2
@@
$\mathcal{E}.1_{T}$ $=$ $\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left|\widehat{\mathbf{c}}_{i}(11)^{-1}_{i}b_{i}(1)}}\right|<\sqrt{T}\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda}{2T}\mathopen{}\mathclose{{\left|\widehat{\mathbf{c}}_{i}(11)^{-1}Sign(\beta_{i}^{*})}}\right|}}\right)}}\right\},$
$i=1,\dots,s,n+1,\dots,n+s_{y}\ \ ,$
$\mathcal{E}.2_{T}$ $=$ $\mathopen{}\mathclose{{\left\{|b_{i}-b_{i}(2)|\leq\frac{\lambda\varphi}{2\sqrt{T}}}}\right\},\ i=s+1,\dots,n,n+s_{y}+1,\dots,n+p_{y}\ \ ,$
where $\widehat{\mathbf{c}}_{i}(11)_{i}$, $b_{i}$, $b_{i}(1)$ and $b_{i}(2)$ are elements of $\widehat{\mathbf{C}}_{w}(11),\ \mathbf{b}=\mathopen{}\mathclose{{\left(\widehat{\mathbf{C}}_{21}(\widehat{\mathbf{C}}_{11})^{-1}\mathbf{W}(1)\mathbf{v}}}\right),\ \mathbf{b}(1)=\frac{1}{\sqrt{T}}\mathbf{W}(1)\mathbf{v}$ and $\mathbf{b}(2)=\sqrt{T}\mathbf{W}(2)\mathbf{v}$, respectively. $\mathcal{E}.1_{T}$ implies that the signs of the relevant predictors are correctly estimated, while $\mathcal{E}.1_{T}$ and $\mathcal{E}.2_{T}$ together imply that the signs of the irrelevant predictors are shrunk to zero. To show $P\mathopen{}\mathclose{{\left(\exists\lambda\geq 0:Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\rightarrow 1$, it is sufficient to show that $P\mathopen{}\mathclose{{\left(\exists\lambda\geq 0:Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\geq P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}\cap\mathcal{E}.2_{T}}}\right)$ (see Proposition 1 in zhaoyu2006). Using the identity of $1-P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}\cap\mathcal{E}.2_{T}}}\right)\leq P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}^{c}}}\right)+P\mathopen{}\mathclose{{\left(\mathcal{E}.2_{T}^{c}}}\right)$ we have that
$P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}^{c}}}\right)+P\mathopen{}\mathclose{{\left(\mathcal{E}.2_{T}^{c}}}\right)\leq\\ \sum_{i=1}^{s,n+1,\dots,n+s_{y}}P\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{T}}|\widehat{\mathbf{c}}_{i}(11)^{-1}\mathbf{w}_{i}^{\prime}\mathbf{v}|\geq\sqrt{T}\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda}{2T}|\widehat{\mathbf{c}}_{i}(11)^{-1}Sign(\beta_{i}^{*})|}}\right)}}\right)\\ +\sum_{i=1}^{s+1,\dots,n,n+s_{y}+1,\dots,n+p_{y}}P\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{T}}|b_{i}-\mathbf{w}_{i}^{\prime}\mathbf{v}|\geq\frac{\lambda\varphi}{2\sqrt{T}}}}\right)=I_{T}+II_{T}\ \ .$ (8)
Note that by Assumption 3 of the main text, $\widehat{\psi}_{max}^{w}\geq\widehat{\psi}_{min}^{w}\geq 0$, hence
@@
for some $c_{0}>0$ (see zhaoyu2006, Thm. 3 and 4). Therefore, by the union bound, Markov’s inequality and the mixingale concentration inequality (see Hansen1991, Lemma 2), we have that
@@
where $c_{1}>2$ (see Assumption 2 ($b$) in the main text). Conducting a similar analysis for $II_{T}$, and considering that by assumption $\sqrt{T}\lambda\rightarrow\lambda_{0}\geq 0$, we obtain $P\mathopen{}\mathclose{{\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\rightarrow 1.$ $\square$
A.5 Proof of Theorem 3
We start by introducing some important definitions.
Definition 1:
@@ A.5 Proof of Theorem 3
Definition 2:
(Mixingale, davidson1994, ch. 16). The sequence of pairs $\mathopen{}\mathclose{{\left\{X_{t},\mathcal{F}}}\right\}_{-\infty}^{+\infty}$ in a filtered probability space $\mathopen{}\mathclose{{\left(\Omega,\mathcal{F},P}}\right)$ where the $X_{t}$ are integrable r.v.s is called $L_{p}$-mixingale if, for $p\geq 1$, there exist sequences of non-negative constants $\mathopen{}\mathclose{{\left\{d_{t}}}\right\}_{-\infty}^{+\infty}$ and $\mathopen{}\mathclose{{\left\{\nu_{m}}}\right\}_{0}^{\infty}$ such that $\nu_{m}\rightarrow 0$ as $m\rightarrow\infty$ and
@@
Definition 3:
(Near-Epoch Dependence, davidson1994, ch. 17). For a possibly vector-valued stochastic sequence $\mathopen{}\mathclose{{\left\{\mathbf{V}_{t}}}\right\}_{-\infty}^{+\infty}$, in a probability space $\mathopen{}\mathclose{{\left(\Omega,\mathcal{F},P}}\right)$ let $\mathcal{F}_{t-m}^{t+m}=\mathbf{\sigma\mathopen{}\mathclose{{\left(\mathbf{V}_{t-m},\dots,\mathbf{V}_{t+m}}}\right)}$, such that $\mathopen{}\mathclose{{\left\{\mathcal{F}_{t-m}^{t+m}}}\right\}_{m=0}^{\infty}$ is a non-decreasing sequence of $\mathbf{\sigma}$-fields. If for $p>0$ a sequence of integrable r.v.s $\mathopen{}\mathclose{{\left\{X_{t}}}\right\}_{-\infty}^{+\infty}$ satisfies
@@
To simplify the analysis, we frequently make use of arbitrary positive finite constants $C$, as well as of its sub-indexed version $C_{i}$, whose values may change from line to line throughout the paper, but they are always independent of the time and cross-sectional dimension. Generic sequences converging to zero as $T\rightarrow\infty$ are denoted by $\zeta_{T}$. We say a sequence $\zeta_{T}$ is of size $\phi_{0}$ if $\zeta_{T}=O(T-\phi_{0}-\varepsilon)$ for some $\varepsilon>0$.
Remark 3:
@@
Under Assumption 1 of the main text the process $\{x_{i,t}\}$ is $L_{2b_{1}}$-NED of size a-a, with $a\geq 1$, while the process $\{q_{i,t}\}$ is $L_{2c_{1}}$-NED of size d-d, with $d\geq 1$. By Theorems 17.5 in ch.17 of davidson1994, they are also $L_{b_{1}}$ and $L_{c_{1}}$-Mixingale, respectively. In later theorems, the NED order and sequence size are important for asymptotic rates. Assumption 2 ($b$) of the main text requires $\mathbf{q}_{t}$ to have slightly more moments than $c_{1}$. More moments mean tighter error bounds and weaker tuning parameter conditions, but a high $c_{2}$ imposes stronger model restrictions. Under strong dependence, fewer moments are needed, and the reduction from $c_{2}$ to $c_{1}$ reflects the cost of allowing greater dependence through a smaller mixing rate.
Proof of Theorem 3 Let $\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}=\sum_{l=1}^{\widehat{p}_{i}}\widehat{\phi}_{i,l}x_{i,t-l}+\sum_{k=1}^{\widehat{q}_{i}}\widehat{\theta}_{i,k}u_{i,t-k}$, $\widehat{\mathbf{x}}_{t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}=(\widehat{x}_{1,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})},\dots,\widehat{x}_{n_{T},t}^{(\boldsymbol{\phi},\boldsymbol{\theta})})^{\prime}$, and $\mathbf{y}_{t-1}^{(p_{y})}=(y_{t-1},\dots,y_{t-p_{y}})^{\prime}$. Note that $\sum_{t=1}^{T}v_{t}\mathbf{w}_{t}^{\prime}=\sum_{t=1}^{T}v_{t}\mathopen{}\mathclose{{\left(\widehat{\mathbf{u}}_{t}^{\prime},\mathbf{y}_{t-1}^{(p_{y})^{\prime}}}}\right)$. Therefore,